Developing a workforce schedule (using Linear Programming to model and solve this problem)

A local bank needs the minimum number of employees needed for each day of the week listed in the following table. If a staff is hired, his/her schedule will be working 5 consecutive days and take two days off. The bank operates seven days a week.

Day of the Week

M

T

W

TH

F

Sa

Su

Number of staff needed

4

5

5

3

5

2

3

So how will you help the bank hire the staffs?
Answer #1

Let,

xi = number of employees start working in day i where i = {Monday=1,Tuesday=2,Wednesday=3,Thursday=4,Friday=5,Saturday=6,Sunday=7}

Objective is to minimize number of employees = Min x1+x2+x3+x4+x5+x6+x7

Subject to,

x1+x4+x5+x6+x7 >=4 (Monday)
x1+x2+x5+x6+x7 >=5 (Tuesday)  
x1+x2+x3+x6+x7 >=5 (Wednesday)  
x1+x2+x3+x4+x7 >=3 (Thursday)  
x1+x2+x3+x4+x5 >=5 (Friday)  
x2+x3+x4+x5+x6 >=2 (Saturday)  
x3+x4+x5+x6+x7 >=3 (Sunday)

xi >= 0 (non-negativity constraint)

Solving in solver we get,

Recommendation to the bank about hiring staff

Minimum number of employee required to hire = 6  

and 3 employee should start working on Monday
1 employee should start working on Wednesday
1 employee should start working on Friday
1 employee should start working on Saturday

Solver screenshot

E А В C G J К L М N 1 2 х1 x2 х3 x4 x5 хб x7 Objective function 3 0 1 0 1 1 0 6 4 5 Monday 6 Tuesday 7 Wednesd 1 1 1 1 1 5 1
Solver formula  

А В C D E F G Н J К L M 1 2 х1 x2 х3 х4 x5 хб x7 Objective function =SUM(B3:H3) 0 1 0 1 1 0 5 Monday 6 Tuesday 7 Wednesday 8